OpenStudy (anonymous):

Find the solution of five times the square root of the quantity of x plus 7 equals negative 10 and determine if it is an extraneous solution. x = −3; extraneous x = −3; not extraneous x = 11; extraneous x = 11; not extraneous

OpenStudy (anonymous):

-3 and extraneous

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Do you need me to explain why it is extraneous?

OpenStudy (anonymous):

that would be nice

OpenStudy (anonymous):

you there or nah?

OpenStudy (anonymous):

\[5\sqrt{x+7}=-10\]\[\sqrt{x+7}=-2\]\[x+7=4\]\[x=-3\] Now if we put x=-3 in the LHS of the original equation we get\[5\sqrt{x+7}=5\sqrt{4}\] But \[\sqrt{4}=2\ and\ NOT\ -2!\] So the LHS=10 and not -10. So the solution is extraneous.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Sorry, typing equations takes a long time.

OpenStudy (anonymous):

its all good thanks man!